OpenStudy (anonymous):

Without graphing, is the system independent, dependent, or inconsistent? y = 2x - 8 2x - y = 8

OpenStudy (anonymous):

both equations are exactly the same.

OpenStudy (anonymous):

2x-y=8 -y= -2x+8 y= 2x-8

OpenStudy (anonymous):

so its independent? or dependent?

OpenStudy (anonymous):

Since both equations are same and there is no added variable hence the equation is independent

OpenStudy (anonymous):

Thank you so much

OpenStudy (phi):

This explains whether dependent, independent or inconsistent http://www.algebra.com/algebra/homework/coordinate/Types-of-systems-inconsistent-dependent-independent.lesson

OpenStudy (anonymous):

@phi or @ipm1988 do you know how to do my other question i posted

OpenStudy (anonymous):

@algebra2sucks post again your question

OpenStudy (anonymous):

----------------------With all due respect, @ipm1988 is incorrect; the answer is dependent for anyone going to connections!! :)-----------------------